# EffLiFe: Efficient Light Field Generation via Hierarchical Sparse Gradient Descent

### v1

@@ Abstract
With the rise of Extended Reality (XR) technology, there is a growing need for real-time light field generation from sparse view inputs. Existing methods can be classified into offline techniques, which can generate high-quality novel views but at the cost of long inference/training time, and online methods, which either lack generalizability or produce unsatisfactory results. However, we have observed that the intrinsic sparse manifold of Multi-plane Images (MPI) enables a significant acceleration of light field generation while maintaining rendering quality. Based on this insight, we introduce EffLiFe, a novel light field optimization method, which leverages the proposed Hierarchical Sparse Gradient Descent (HSGD) to produce high-quality light fields from sparse view images in real time. Technically, the coarse MPI of a scene is first generated using a 3D CNN, and it is further sparsely optimized by focusing only on important MPI gradients in a few iterations. Nevertheless, relying solely on optimization can lead to artifacts at occlusion boundaries. Therefore, we propose an occlusion-aware iterative refinement module that removes visual artifacts in occluded regions by iteratively filtering the input. Extensive experiments demonstrate that our method achieves comparable visual quality while being 100x faster on average than state-of-the-art offline methods and delivering better performance (about 2 dB higher in PSNR) compared to other online approaches.
@@
Light field, Multi-plane Image, Hierarchical Sparse Gradient Descent, Occlusion-aware Iterative Refinement.
@@ 1 Introduction
Figure 1: Derivation of MPI sparse gradient and optimization process. (a) An MPI plane $M_{d}$ at depth $d$ can be decomposed into an RGB plane $c_{d}$ and a transparency/alpha plane $\alpha_{d}$. (b) The multi-plane alpha gradients $\mathcal{A}_{d}$ are derived from multi-plane alpha planes by calculating the gradient of $M$ corresponding to $c_{d}$, and $\mathcal{A}_{d}$ is then sparsified by clipping small gradients, which is used to refine the MPI using our proposed sparse gradient descent strategy.
Figure 2: Application of our method to support a real-time 3D display. (a) Sparse multi-view images that serve as the input to our model. (b) Our Hierarchical Sparse Gradient Descent (HSGD) is capable of generating multi-plane images (c) online at around 30 FPS. (d) Novel views can then be rendered offline from a Multi-plane Image at approximately 700 FPS. (e) Several novel views rendered from the MPI provide enough light field information to a 3D display, specifically a looking glass that supports naked-eye 3D effects from a wide range of views. (GD is short for gradient descent.)
@@
With the rapid development of XR/Naked Eye 3D display devices, there is an increasing demand for live light field video rendering technique that can record and stream light field video in real time, allowing viewers to interactively change the viewpoint and focus of the video. Even with the latest breakthroughs of light field generation[1, 2] and neural radiance fields[3, 4], it still remains a challenging task.
@@
To enable live light field generation, the supporting algorithms and representations should be capable of producing high-quality view synthesis results in real time and generalize well to unseen scenes. However, current approaches face difficulties in achieving a balance between visual performance and real-time effectiveness:
@@
  • Methods that rely on implicit scene representation, such as NeRF [3] and its derivatives [5, 6, 7, 8] can produce view synthesis results with high visual quality, but require long-time per-scene optimization and dense view inputs. While some generalization techniques [9, 10, 11, 12, 13, 14] can generate novel views that are generalizable with the help of multi-view stereo (MVS) geometry rules, it is still not fast enough for them to achieve real-time performance due to high memory usage and complex computation of intermediate cost volumes.
@@
  • Methods that rely on explicit scene representation, such as Plenoxels[15], PlenOctrees[16] and DVGO[17] enclose a 3D scene in a voxel grid and optimize the radiance fields within it. While such methods can reproduce intricate geometric and texture details at high grid resolutions, they can not generalize to new scenes without extensive per-scene optimization. While some methods with learned features [18, 19, 20, 1] can generalize well to unseen scenes from sparse views, they still require a significant amount of time to infer a scene due to the heavy computational burden of processing large-sized MPIs and plane sweep volumes.
@@
  • Methods relying on surrogate geometry are computationally efficient by only sampling around the estimated geometry scaffolds such as depth maps [21] and surface mesh. However, these methods may fail if the precomputed geometry scaffolds are incorrectly estimated for semi-transparent surfaces or thin plates, therefore they are not very robust for complex scenarios.
Based on the above observations, we conclude that live light field reconstruction remains a challenging task for the following reasons: 1) Recovering a 3D light field from 2D images is a difficult ill-posed inverse problem that necessitates either dense input views or geometric priors learned from data; 2) Implicit representations save memory but are computationally complex, resulting in a long inference time. Explicit representations offer faster rendering speed but require significant memory and time for constructing the light field; 3) High-quality light field generation necessitates sophisticated optimization to handle intricate geometry details such as semi-transparent surfaces and thin structures.
To overcome the challenges, we propose EffLiFe, which employs Hierarchical Sparse Gradient Descent (HSGD) to produce high-quality light fields in real time from sparsely sampled views. Unlike prior methods[18, 19, 1] that directly generate a full-sized MPI, our approach derives it hierarchically, as light field details within an MPI can be effectively refined utilizing high-resolution information grounded on a well-established manifold. Traditionally, an MPI is estimated by minimizing disagreement between predicted and observed measurements using analytical gradient descent. DeepView[1] incorporates a learned gradient descent strategy[22, 23] that explicitly computes gradients of MPI’s color and alpha channels. These gradients are fed into a neural network for rapid and adaptive gradient descent. However, in contrast to DeepView [1] that utilizes full-sized MPI gradients, our approach selectively employs important gradients, based on the observation that useful gradients in the MPI updating are spatially sparse. Specifically, inverse rendering utilizes gradient descent to refine the desired scene representation, while the majority of gradients are close to zero. This implies that further optimization in these areas is unlikely to enhance rendering quality and would instead demand significant computational resources.
@@
Fig. 1 depicts the derivation of our MPI sparse alpha gradients which are utilized for sparse gradient descent. Firstly, the multi-plane gradients $\mathcal{A}_{d}$ are derived by computing the gradients of the MPI with respect to its color channel. Next, $\mathcal{A}_{d}$ is sparsified by clipping small gradients. Finally, the sparse gradient descent module updates the MPI according to the sparse important gradients. Experiments on public datasets demonstrate that compared with previous methods[18, 19, 1], our optimization strategy based on the sparsity of gradients significantly reduces the computational complexity without loss of accuracy. Also, in comparison with methods[24] relying on surrogate geometry, our method achieves superior visual performance while maintaining high temporal efficiency.
Typically, novel view synthesis approaches suffer from poor quality at occlusion boundaries, which can result in unpleasant visual artifacts. Although these artifacts may have a limited impact on standard metrics such as PSNR/SSIM, they can still be noticeable and negatively impact the overall viewing experience. Thus, we further propose an occlusion-aware iterative refinement module to address this issue without introducing additional complexity. We identify that ambiguity at occlusion boundaries is mainly caused by inconsistent observations across input views. More specifically, an MPI is accurately estimated only if observations from all input views correspond well to each other, but occluded views provide random noisy observations making the ill-posed inverse problem hard to converge. Based on this observation, we identify the occluded regions based on the previously estimated scene representation and replace them with observations from other non-occluded views.
@@
  • We present an optimization strategy: Hierarchical Sparse Gradient Descent (HSGD), that efficiently recovers high-quality light fields while effectively reducing inference time and GPU memory usage.
  • We propose an occlusion-aware iterative refinement module that filters the input to reduce occlusion artifacts. This method is free of learning and performs effectively for scenes with severe occlusion.
  • Extensive experiments on public datasets demonstrate that EffLiFe achieves a PSNR improvement of approximately 2db compared to other online approaches, or comparable performance while being on average 100x faster than offline approaches.
  • As a 3D display application demonstrated in Fig. 2, our solution can generate the MPIs at around 30 FPS, and novel views can be rendered from a built-up MPI at about 700 FPS for the resolution of $378\times 512$. Novel views of different angles are then provided to a 3D display, which presents naked-eye 3D effects through multi-layer light field decomposition and directional backlighting [25].
Figure 3: The overview of EffLiFe. First, the plane sweep volume (PSV) is constructed using multi-view images by homographic warping, and the PSV is then downsampled hierarchically at multiple resolutions. Thereafter, the output MPI is generated in several iterations. Initially, the lowest-resolution PSV is fed to a 3D CNN to extract a coarse-level multi-plane image (MPI). For the subsequent iteration, occlusion of the PSV is first filtered guided by the upsampled MPI of the previous level. The filtered PSV and the upsampled MPI both go through the Sparse Gradient Descent module for a refined higher-resolution MPI. This loop is repeated until the resolution of the output MPI reaches the same resolution with the input images. Finally, a novel view can be easily rendered from the MPI using the repeated over operator[26].
@@ 2 Related Work
We focus on live light field generation in this paper, which requires novel view synthesis that generalizes to arbitrary scenes with high rendering quality and can be processed in real time. However, existing approaches can hardly achieve all these requirements.
@@ 2.1 High-quality Novel View Synthesis
The rendering quality of novel view synthesis has seen significant improvement since the proposal of NeRF[3]. NeRF encodes a scene in a 3D radiance volume and optimizes the density and color of the volume using a simple MLP. Novel views are generated by integrating the volume’s density and color along each viewing ray. Building upon the success of NeRF, several methods such as Mip-NeRF[4], RawNeRF[27], Ref-NeRF[28], and Mip-NeRF360[29] have further improved the rendering quality of novel view synthesis. For example, Mip-NeRF reduces objectionable aliasing artifacts by casting cones and prefiltering the positional encoding function, while RawNeRF trains the NeRF model on raw camera data to achieve novel high dynamic range view synthesis. Ref-NeRF produces realistic shiny surfaces by explicitly modeling surface reflections, and Mip-NeRF360 solves the problem of unbounded scene novel view synthesis with a non-linear scene parameterization, online distillation, and a novel regularizer. Additionally, another MPI-based method[18] parameterizes each pixel as a linear combination of basis neural functions, successfully reproducing realistic view-dependent effects.
@@ 2.3 Real-time Novel View Synthesis
To achieve real-time inference speed of a trained scene, structured representations and efficient computation skills are employed. KiloNeRF[7] replaces the original large MLP of NeRF with thousands of smaller faster-to-evaluate MLPs, enabling up to 40 FPS of rendering speed. Plenoctrees[16] render novel views by representing the scene in a structured octree-based 3D grid. FastNeRF[8] is capable of rendering novel views at 200 FPS by caching a deep radiance map and efficiently querying it. Except for NeRF-like representations, a built-up multi-plane image (MPI) from StereoMag/LLFF/DeepView[18, 19, 1] can achieve a rendering speed of 60 FPS. (The rendering speeds of above methods are all evaluated on 800x800 images of Synthetic-NeRF[3] dataset.) Despite the real-time inference speed of novel view synthesis, real-time building-up of a new scene is hard to be reached. The cutting-edge optimization-based method: InstantNGP[32] reduces training time of NeRF from hours to several seconds by using a hash grid and hardware acceleration techniques. DeepView[1] is now the fastest MPI-based method, however, it still requires about 50 seconds to generate an MPI. ENeRF[24] is the state-of-the-art generalizable novel view synthesis method and can render new views at 30 FPS on the ZJU-MoCap[33] dataset by sampling few points around a computed depth map, however, the rendering speed and image quality can still be improved.
2.4 Light Field Video Streaming
Instead of producing a single image at a time, [18, 19, 1, 2, 34, 30] output the entire light field of the scene as multi-plane images or layered mesh representations. With a generated scene light field, novel views can be rendered easily with nearly no computation budget by simply ray querying. [2] encodes the scene in a set of multi-sphere images and compresses them into an atlas of scene geometry, allowing light field videos to be streamed over a gigabit connection. Though this method is real-time in rendering and streaming, it still requires a long time to generate a new scene light field because it is based on DeepView[1]. Starline[35] and VirtualCube[36] are another two methods that make live light field streaming possible. They both build up their systems using several RGBD cameras, and the additional depth channel enables fast scene geometry extraction, thus saving much time for computation. However, in this paper, we only compare RGB-based novel view synthesis methods to explore their potential for live light field generation.
@@ 3 Method
The proposed method, EffLiFe, aims to facilitate real-time generation of light fields using a collection of posed source view images. To accomplish this goal, we introduce Hierarchical Sparse Gradient Descent (HSGD), an efficient and high-quality optimization technique for the Multi-plane Image. By leveraging HSGD, we can achieve real-time light field generation while maintaining a high level of quality.
Fig. 3 presents an overview of the proposed method, which comprises two primary processes: data preparation and hierarchical sparse gradient descent. Sec. 3.2 introduces the data preparation stage, which encompasses the construction of the plane sweep volume and the initial generation of the Multi-plane Image (MPI). Sec 3.3 introduces the hierarchical sparse gradient descent stage, as depicted in Fig. 5. This module comprises three major operations: gradient formulation, sparse gradient selection, and sparse gradient update. Prior to discussing our method in detail, we provide a concise introduction of the Multi-plane Image (MPI), and our fundamental optimization method, Learned Gradient Descent, in Sec. 3.1.
Figure 4: The sparsity of MPI gradients. (a) The rendered images of two MPIs. (b) Randomly selected two multi-plane gradients corresponding to the color channel of two MPIs.
Figure 5: The sparse gradient descent module. (a) The combined volume $V$ comprises the input plane sweep volume $P$, the warped alpha gradients $\mathcal{A}_{i}^{w}$ for each source view $i$, and the input multi-plane image $M_{r}$. (b) The volume $V$ is sparsified by selecting the top $k$ pixels along the depth axis based on alpha gradients. Simultaneously, the output sparse indices $S$ store the positions of the selected $k$ pixels. (c) The sparsified volume $V_{s}$ is fed to a 3D CNN for a refiend sparse MPI gradient. Finally, the sparse gradient update module utilizes the sparse indices $S$ to add the sparse MPI gradient to the input multi-plane image $M_{r}$, resulting in a refined multi-plane image.

### v3

# RealLiFe: Real-Time Light Field Reconstruction via Hierarchical Sparse Gradient Descent

@@ Abstract
With the rise of Extended Reality (XR) technology, there is a growing need for real-time light field reconstruction from sparse view inputs. Existing methods can be classified into offline techniques, which can generate high-quality novel views but at the cost of long inference/training time, and online methods, which either lack generalizability or produce unsatisfactory results. However, we have observed that the intrinsic sparse manifold of Multi-plane Images (MPI) enables a significant acceleration of light field reconstruction while maintaining rendering quality. Based on this insight, we introduce RealLiFe, a novel light field optimization method, which leverages the proposed Hierarchical Sparse Gradient Descent (HSGD) to produce high-quality light fields from sparse input images in real time. Technically, the coarse MPI of a scene is first generated using a 3D CNN, and it is further optimized leveraging only the scene content aligned sparse MPI gradients in a few iterations. Extensive experiments demonstrate that our method achieves comparable visual quality while being 100x faster on average than state-of-the-art offline methods and delivers better performance (about 2 dB higher in PSNR) compared to other online approaches.
@@
Light field, Multi-plane Image, Hierarchical Sparse Gradient Descent.
@@ 1 Introduction
With the rapid development of XR / naked eye 3D display devices, there is an increasing demand for live light field video rendering techniques that can record and stream light field video in real time, allowing viewers to interactively change the viewpoint and focus of the video. Even with the latest breakthroughs in light field reconstruction[1, 2] and neural radiance fields[3, 4], it remains a challenging task.
Figure 1: Rendering quality and efficiency comparison with state-of-the-art novel view synthesis methods[5, 6, 7] and light field reconstruction methods[8] on Real Forward-Facing[9] of image size $512\times 384$. RealLiFe is our default model with 3 iterations of gradient descent, and RealLiFe-2I is one with 2 iterations of gradient descent.
Figure 2: The derivation of MPI sparse gradients and the optimization process. (a) An MPI plane $M_{d}$ at depth $d$ can be decomposed into an RGB plane $c_{d}$ and an alpha (transparency) plane $\alpha_{d}$. (b) The multi-plane alpha gradients $\mathcal{A}_{d}$ are computed from multi-plane alpha planes by calculating the gradient of $M$ corresponding to $c_{d}$, and $\mathcal{A}_{d}$ is then sparsified by discarding small gradients, which is used to refine the MPI using our proposed sparse gradient descent strategy.
Figure 3: Application of our method to support a real-time 3D display. (a) Sparse multi-view images that serve as the input to our model. (b) Our Hierarchical Sparse Gradient Descent (HSGD) is capable of generating multi-plane images (c) online at around 35 FPS. (d) Novel views can then be rendered offline from a Multi-plane Image at approximately 700 FPS. (e) Several novel views rendered from the MPI provide enough light field information to a 3D display, specifically a looking glass that supports naked-eye 3D effects from a wide range of views. (GD is short for gradient descent.)
@@
To enable live light field reconstruction, the supporting algorithms and representations should be capable of producing high-quality view synthesis results in real time and generalize well to unseen scenes. However, current approaches face difficulties in achieving a balance between visual performance and real-time effectiveness.
@@
  • Methods relying on implicit scene representation, such as NeRF [3] and its derivatives [10, 11, 12, 13] can produce view synthesis results with high visual quality, but require long-time per-scene optimization and dense view inputs. While some generalization techniques [14, 15, 16, 5, 6, 17] can generate novel views that are generalizable with the help of multi-view stereo (MVS) geometry rules, it is still not fast enough for them to achieve real-time performance due to high memory usage and complex computation of intermediate cost volumes.
@@
  • Methods that rely on explicit scene representation, such as Plenoxels[18], PlenOctrees[19] and DVGO[20] enclose a 3D scene in a voxel grid and optimize the radiance fields within it. While such methods can reproduce intricate geometric and texture details at high grid resolutions, they can not generalize to new scenes without extensive per-scene optimization. Although some methods with learned features [21, 9, 22, 1] can generalize well to unseen scenes from sparse views, they still require a significant amount of time to infer a scene due to the heavy computational burden of processing large multi-plane image (MPI) and plane sweep volume (PSV).
@@
  • Methods relying on surrogate geometry are computationally efficient by sampling only around the estimated geometry scaffolds, such as depth maps [23] and surface mesh. However, these methods may fail if the precomputed geometry scaffolds are incorrectly estimated for semi-transparent surfaces or thin plates; therefore, they are not very robust for complex scenarios.
Based on the above observations, we conclude that live light field reconstruction remains a challenging task for the following reasons:
1) Recovering a 3D light field from 2D images is a difficult ill-posed inverse problem that necessitates either dense input views or geometric priors learned from data;
2) Implicit representations save memory but are computationally complex, resulting in a long inference time. Explicit representations offer faster rendering speed but require significant memory and time for constructing the light field;
3) High-quality light field reconstruction necessitates sophisticated optimization to handle intricate geometry details such as semi-transparent surfaces and thin structures.
To overcome the challenges, we propose RealLiFe, a novel method that employs Hierarchical Sparse Gradient Descent (HSGD) to produce high-quality light fields in real time from sparsely sampled views.
Unlike previous methods[21, 9, 1] that directly process and generate a full MPI, our method adopts a hierarchical approach, progressively refining it from low resolution to high resolution. This approach draws inspiration from the principles of Multi-View Stereo (MVS), where intricate geometric details can be iteratively enhanced with high-resolution features and well-established low-resolution geometric structures[24, 25, 26, 27, 28, 29]. We extend this concept from geometry to appearance, claiming and proving that the visual details within a light field can also be iteratively improved, given access to high-resolution captures and well-structured low-resolution representations.
[…]
Before the design for efficiency, it is imperative to address the fundamental optimization process for deriving an MPI from sparse views. This task poses a significant challenge, as it involves solving an inverse problem. Conventionally, this is accomplished by iteratively minimizing the discrepancies between the predicted and observed measurements using analytic gradient descent, which can be a time-consuming process. However, DeepView[1] introduces a novel approach by incorporating the learned gradient descent strategy[30, 31]. This strategy enables the rapid generation of an MPI in only a few iterations by explicitly computing MPI gradients and relying on the neural network to perform adaptive gradient descent. We also apply this strategy for rapid gradient descent, but, in contrast to DeepView [1], which utilizes full-sized and full-channel MPI gradients, our approach only employs sparsely gradients that align with the scene contents. This choice arises from the observation that gradients within vacuous regions either insignificantly affect or negligibly contribute to the final rendering quality.
Based on our observations, we present the fundamental stages of the proposed Hierarchical Sparse Gradient Descent (HSGD), as depicted in Fig. 2. For example, in one stage, the multi-plane gradients $\mathcal{A}_{d}$ are initially obtained by computing the gradients of the MPI with respect to its RGB channel, as outlined in DeepView[1]. Subsequently, $\mathcal{A}_{d}$ undergoes sparsification by removing small-value gradients. Finally, the sparse gradient descent module updates the MPI based on the sparse gradients.
@@
Experiments on public datasets demonstrate that compared to previous methods[21, 9, 1], our optimization strategy based on the sparsity of gradients significantly reduces computational complexity without loss of accuracy. Also, in comparison with methods[7] relying on surrogate geometry, our method achieves superior visual performance while maintaining high temporal efficiency.
@@
  • We present an optimization strategy: Hierarchical Sparse Gradient Descent (HSGD), which efficiently recovers high-quality light fields while effectively reducing inference time and GPU memory usage.
  • Extensive experiments on public datasets demonstrate that RealLiFe achieves a PSNR improvement of approximately 2db compared to other online approaches, or comparable performance while being on average 100x faster than offline approaches.
  • As a 3D display application demonstrated in Fig. 3, our solution can generate MPIs at above 35 FPS, and novel views can be rendered from a built-up MPI at about 700 FPS for the resolution of $378\times 512$. Novel views of different angles are then provided to a 3D display, which presents 3D effects in the naked eye through multilayer light field decomposition and directional backlighting [32].
Figure 4: The overview of RealLiFe. First, the PSV is constructed using multiview images by homographic warping, and the PSV is then downsampled hierarchically at multiple resolutions. The output MPI is then generated in several iterations. Initially, the lowest-resolution PSV is fed to a 3D CNN to extract a coarse-level MPI. The PSV and the upsampled MPI both go through the Sparse Gradient Descent module for a refined higher-resolution MPI. This loop is repeated until the resolution of the output MPI reaches the same resolution as that of the input images. Finally, a novel view can be easily rendered from the MPI using the repeated over operator[33].
@@ 2 Related Work
We focus on live light field reconstruction in this paper, which requires novel view synthesis that generalizes to arbitrary scenes with high rendering quality and can be processed in real time. However, existing approaches can hardly achieve all these requirements.
@@ 2.1 High-quality Novel View Synthesis
The rendering quality of novel view synthesis has seen significant improvement since the proposal of NeRF[3]. NeRF encodes a scene in a 3D radiance volume and optimizes the density and color of the volume using a simple MLP. Novel views are generated by integrating the volume’s density and color along each viewing ray. Based on the success of NeRF, several methods such as Mip-NeRF[4], RawNeRF[34], Ref-NeRF[35], and Mip-NeRF360[36] have further improved the rendering quality of the novel view synthesis. For example, Mip-NeRF reduces objectionable aliasing artifacts by casting cones and prefiltering the positional encoding function, while RawNeRF trains the NeRF model on raw camera data to achieve novel high dynamic range view synthesis. Ref-NeRF produces realistic shiny surfaces by explicitly modeling surface reflections, and Mip-NeRF360 solves the problem of unbounded scene novel view synthesis with a non-linear scene parameterization, online distillation, and a novel regularizer. Additionally, another MPI-based method[21] parameterizes each pixel as a linear combination of basis neural functions, successfully reproducing realistic view-dependent effects.
@@ 2.3 Real-time Novel View Synthesis
To achieve real-time inference speed of a trained scene, structured representations and efficient computation skills are employed. KiloNeRF[12] replaces the original large MLP of NeRF with thousands of smaller faster-to-evaluate MLPs, enabling up to 40 FPS of rendering speed. Plenoctrees[19] render novel views by representing the scene in a structured octree-based 3D grid. FastNeRF[13] is capable of rendering novel views at 200 FPS by caching a deep radiance map and efficiently querying it. Except for NeRF-like representations, a built-up MPI from StereoMag/LLFF/DeepView[21, 9, 1] can achieve a rendering speed of 60 FPS. (The rendering speeds of above methods are all evaluated on 800x800 images of Synthetic-NeRF[3] dataset.) Despite the real-time inference speed of novel view synthesis, real-time building-up of a new scene is hard to be reached. The cutting-edge optimization-based method: InstantNGP[38] reduces training time of NeRF from hours to several seconds by using a hash grid and hardware acceleration techniques. DeepView[1] is now the fastest MPI-based method, however, it still requires about 50 seconds to generate an MPI. ENeRF[7] is the state-of-the-art generalizable novel view synthesis method and can render new views at 30 FPS on the ZJU-MoCap[39] dataset by sampling few points around a computed depth map; however, the rendering speed and image quality can still be improved.
2.4 Light Field Reconstruction
Instead of producing a single image at a time, there are methods that output the entire light field of the scene as multi-plane images[21, 9, 1] or layered mesh representations[2, 40, 8]. With a generated scene light field, novel views can be rendered easily with nearly no computation budget by simply ray querying. [2] encodes the scene in a set of multi-sphere images and compresses them into an atlas of scene geometry, allowing light field videos to be streamed over a gigabit connection. Though this method is real-time in rendering and streaming, it still requires a long time to reconstruct a new scene light field. [41, 42, 43] optimized for specific neural architectures for improved reconstruction quality of light fields. Their sophisticated designs for the light field capture systems and neural networks result in high metrics in rendering quality but also make them offline light field reconstruction methods. Starline[44] and VirtualCube[45] are another two methods that make live light field streaming possible. They both build up their systems using several RGBD cameras, and the additional depth channel enables fast scene geometry extraction, thus saving much time for computation. However, in this paper, we only compare RGB-based novel view synthesis methods to explore their potential for live light field reconstruction.
@@ 3 Method
The proposed method, RealLiFe, aims to facilitate real-time reconstruction of light fields using a set of sparse posed source view images. To accomplish this goal, we introduce Hierarchical Sparse Gradient Descent (HSGD), an efficient and high-quality optimization technique for MPIs.
Fig. 4 presents an overview of the proposed method, which comprises two primary processes: initial MPI generation and hierarchical sparse gradient descent. Sec. 3.2 introduces the initial MPI generation stage, which encompasses the construction of the plane sweep volume and the initial generation of the MPI. Sec. 3.3 introduces the hierarchical sparse gradient descent stage, as depicted in Fig. 5. This module comprises three major operations: gradient formulation, gradient sparsification, and sparse gradient update. Prior to discussing our method in detail, we provide a concise introduction of the MPI, and our fundamental optimization method, Learned Gradient Descent, in Sec. 3.1. Finally, we provide an introduction of our training approach, with a primary focus on how we construct the loss function, as outlined in Sec. 3.4.
Figure 5: The sparse gradient descent module. (a) The MPI gradients $\nabla$ comprises the input plane sweep volume $P$ and the warped alpha gradients $\mathcal{A}_{i}^{w}$ for each source view $i$. (b) The volume $V$ is composed of the MPI $M$ and the MPI gradients $\nabla$. It is sparsified by selecting the top $k$ voxels along the depth axis based on alpha gradients of the reference view. Simultaneously, the output sparse indices $S$ store the positions of the selected $k$ voxels along the depth axis. (c) The sparsified volume $V_{s}$ is fed to a 3D CNN (learned gradient descent) for a refiend sparse MPI residual. Finally, the sparse gradient update module utilizes the sparse indices $S$ to add the sparse MPI residual to the input multi-plane image $M$, resulting in a refined multi-plane image.